# Tac-Man: Tactile-Informed Prior-Free Manipulation of Articulated Objects

### v1

@@ III-C Contact Representation
For the effective computation of Eq. 6 and optimization of Eq. 7, a suitable representation of $C_{i}$ is crucial. To ensure compatibility with the homogeneous transformation matrix and to match the capabilities of current tactile sensors, we represent the contact $C_{i}$ as a set of positions of all discretized points at the contact site, relative to the reference frame $\{i\}$. The position of each contact point within $C_{i}$ is denoted by $\mathbf{p}^{i}$. To facilitate alignment with the homogeneous transformation matrix, the column vector representation of $\mathbf{p}^{i}$ is defined in an augmented format as follows:
$\mathbf{p}^{i}=\begin{bmatrix}p_{x}^{i}\\ p_{y}^{i}\\ p_{z}^{i}\\ 1\end{bmatrix},$ (8)
@@
Significantly, $C_{0}$ includes all points on the gripper’s outer surface. Assuming the total number of points is $m$, $C_{0}$ is represented as:
$C_{0}=\{\mathbf{p}_{i}^{0}\mid i=1,\dots,m\}.$ (9)
For simplicity, we assume a flat gripper surface, a common assumption in current robotic grippers and easily achievable with tactile sensor fabrication. Consequently, $p_{x}^{0}$ is consistent across all points and set to 0, attainable through coordinate transformation. The position of a contact point at $C_{i}$ is then characterized by the condition:
$|\mathbf{p}^{i}_{[1]}|\geq\epsilon$ (10)
@@
Therefore, $C_{i}$ for $i=1\dots n$ can be defined as:
$C_{i}=\{\mathbf{p}^{i}\mid|\mathbf{p}^{i}_{[1]}|\geq\epsilon\}.$ (11)
III-D Stable Contact
@@
The computation of stable contact involves addressing the gripper’s maximum elastic deformation, encompassing both normal and shear components. These are captured in the expressions for $f_{d}(C_{i},C_{0})$ and $D$ within the context of stable contact. Ensuring that the contact deformation remains within the material’s elastic limits is crucial for maintaining a stable grip on the object:
$\small f_{d}(C_{i},C_{0})=\begin{bmatrix}\max\{|\mathbf{u}_{[1]}|\mid\mathbf{u}\in C_{i}\}\\ \max\{\|\begin{bmatrix}\mathbf{u}_{[2]}-\mathbf{v}_{[2]}\\ \mathbf{u}_{[3]}-\mathbf{v}_{[3]}\end{bmatrix}\|_{2}\mid(\mathbf{u},\mathbf{v})\in\mathcal{K}_{0i}\}\end{bmatrix}$ (12)
@@
$D=\begin{bmatrix}d_{n}\\ d_{s}\end{bmatrix},$ (13)
where $d_{n}$ and $d_{s}$ denote the maximum normal and shear elastic deformation, respectively. For notation simplicity, we denote $|\mathbf{u}_{[1]}|$ as $\Delta N$ and $\|\begin{bmatrix}\mathbf{u}_{[2]}-\mathbf{v}_{[2]}\\ \mathbf{u}_{[3]}-\mathbf{v}_{[3]}\end{bmatrix}\|_{2}$ as $\Delta S$ henceforth.
Similarly, we formulate $f_{s}(C_{i},C_{0})$ and $S$ to address concerns related to slipping in Eq. 2. Practically, to prevent sliding, the friction $f$ between the object and the gripper should not surpass the maximum static friction $f_{M}$. In our model, the friction at $C_{i}$ is positively correlated with the shear deformation observed from $C_{0}$:
@@
$\forall(\mathbf{u},\mathbf{v})\in\mathcal{K}_{0i},\Delta S\leq\delta^{M}(f_{N}),$ (15)
@@
$\delta^{M}(\Delta S)=\delta_{0}\Delta S,(\mathbf{u},\mathbf{v})\in\mathcal{K}_{0i},$ (16)
@@
$\delta_{0}=\frac{\delta^{M}(\Delta S_{0})}{\Delta S_{0}},(\mathbf{u},\mathbf{v})\in\mathcal{K}_{0i}.$ (17)
@@
$\forall(\mathbf{u},\mathbf{v})\in\mathcal{K}_{0i},\frac{\Delta S}{\Delta N}\leq\delta_{0},$ (18)
leading to the expression for $f_{s}(C_{i},C_{0})$:
$f_{s}(C_{i},C_{0})=\max\{\frac{\Delta S}{\Delta N}\mid(\mathbf{u},\mathbf{v})\in\mathcal{K}_{0i}\},$ (19)
@@
$f_{e}(C_{i},C_{1})=\frac{\sum_{(\mathbf{u},\mathbf{v})\in\mathcal{K}_{1i}}\|\mathbf{u}-\mathbf{v}\|_{2}}{\#\mathcal{K}_{1i}}.$ (21)
@@ III-E Robot Pose Update
Having established the contact representation and computation methods for stable contact constraints, we should address the optimization in Eq. 7 to compute the value of $y$ for the determination of $T_{\mathcal{D}}$. However, in the absence of knowledge about the handle’s trajectory $\mathcal{J}$, determining $y$ explicitly remains challenging. We, therefore, adopt an exploration mechanism by continuously executing with small increments in $y$ until any term in the constraints of optimization Eq. 7 reaches its upper bound, scaled by a safety margin $\alpha$ ($10<\alpha<$ 1). This process is concluded as
$T^{\prime}_{i}=T_{i}T_{\mathcal{D}},$ (22)
@@
To recover the contact to within the specified bounds after the execution stage, we employ an optimization algorithm [70] to compute the optimal transformation $T_{R_{i}}$ for the recovery term in Eq. 6:
$T_{R_{i}}=\operatorname*{arg\,min}_{T_{R_{i}}\in\mathbb{R}^{4\times 4}}\sum_{(\mathbf{u},\mathbf{v})\in\mathcal{K}_{1i}}\|T_{i+1}^{i}\mathbf{u}-\mathbf{v}\|_{2}.$ (23)
@@
Figure 2: Real-world experiment setup. The experiments are conducted in a real-world environment using a Gen3 7DoF robotic arm paired with a Robotiq 2F-85 gripper. Integral to the setup are two GelSight-inspired tactile sensors, each equipped with a $7\times 7$ grid of labeled markers to enhance tactile sensing capabilities.
@@ IV-A Experimental Setup
Figure 3: Visualization of tactile sensor reading. The GelSight-inspired sensors are adept at providing detailed contact information during object interactions. These sensors excel in capturing $\Delta N$ in high spatial resolution, essential for precise detection of contact points. To further enhance their effectiveness, labeled markers are used for accurate point correspondence. Utilizing these features, $\Delta S$ is calculated and visualized using white arrows in the illustration. For better visibility, the magnitude of these arrows is exaggerated by a factor of $5$, ensuring a clearer visualization of the contact dynamics.
@@
Tac-Man is designed for broad compatibility with robotic systems capable of facilitating 6D movement in the gripper. In experiments, we utilized a Kinova Gen3 7DoF robotic arm equipped with a Robotiq 2F-85 gripper (Fig. 2). To assess the contact constraints as outlined in Sec. III, we integrated two GelSight-inspired tactile sensors into the gripper.
GelSight-inspired tactile sensor
Further enhancing our ability to validate these constraints, we based our sensors on the standard GelSight design, which provides high-resolution, pixel-level data at the contact surface. Each sensor was specially fabricated with a unique $7\times 7$ grid of labeled markers (Fig. 3). This grid aids in accurately identifying point correspondences between two contacts. Although this modification slightly reduces the resolution and depth reconstruction accuracy as compared to conventional GelSight sensors, its effectiveness in practical scenarios and in validating the contact constraints detailed in Sec. III is notable.
Figure 4: Selection concerns for the $\epsilon$ value. The flexibility inherent in the tactile sensor material results in the activation of both the direct contact region and the adjacent transitional areas when contacting the handle. This characteristic underlines the need for a carefully adaptable $\epsilon$ value. Such adaptability ensures the precise identification of true contact points, thereby reducing the likelihood of detecting unreliable points.
@@
  • The threshold for identifying contact points, $\epsilon$ in Eq. 10, is set to be adaptable. This is due to the material’s flexibility, as shown in Fig. 4, which results in deformation not only in the contact region but also in the surrounding area upon contact with the object. This deformation relates to $\Delta N$ and the handle’s shape, among other factors. For adaptability, $\epsilon$ is defined as the minimal $\Delta N$ that ensures at least 8 contact points, exceeding the necessary 6 for accurate 6D transformation solving in Eq. 23 and surpassing the number of markers in a line (7 in our sensors), thus ensuring a 3D spread of contact points.
  • The normal elastic deformation limit, $d_{n}$ in Eq. 13, is set at $2~{}\mathrm{mm}$. This value is chosen based on observations that within this range, the system can revert to its initial state after releasing contact. It is important to note that this value represents a conservatively reduced normal elastic deformation for enforcing stricter constraints.
@@
  • The shear elastic deformation limit, $d_{s}$ in Eq. 13, is calculated as $d_{n}\delta_{0}$. Though smaller than the actual one, this value is deemed acceptable for stricter constraint enforcement.
  • The upper bound for another stable contact constraint, $E$ in Eq. 3, is experimentally determined to be $0.4~{}\mathrm{mm}$, ensuring it does not exceed $d_{n}$, $d_{s}$, and $\delta_{0}$ during manipulation.
  • The movement magnitude in the execution stage, $y$ in Eq. 7, is set as $\epsilon\delta_{0}$ to preserve stable contact during execution.
@@
hyper-parameter $\epsilon$ $d_{n}$ $d_{s}$ $\delta_{0}$ $E$ y $\alpha$
value adaptive $2~{}\mathrm{mm}$ $3~{}\mathrm{mm}$ $1.5$ $0.4~{}\mathrm{mm}$ $\epsilon\delta_{0}$ $0.6$
@@
Figure 5: $\delta_{0}$ value acquisition. The experiment to obtain hyper-parameter $\delta_{0}$, as defined in Eq. 17: (a) The experiment begins with the arm pulling backward after the gripper has firmly grasped a slippery Lego piece. (b) Multiple trials are conducted at different values of $\Delta N$. Three representative trials are visualized to illustrate the process. The results follow the friction model well and $\delta^{M}$ can be easily identified. The results align well with the friction model, allowing for the straightforward identification of $\delta^{M}$. (c) The $\delta_{0}$ obtained from these multiple trials is visualized, and a value of $1.5$ is selected for safety considerations.
@@ IV-B Manipulation under Ambiguous Priors
Figure 6: Manipulation under ambiguous priors. Tac-Man showcases its proficiency in addressing the challenge caused by ambiguous priors, characterized by objects that, while visually indistinguishable, feature distinct articulation mechanisms. Our evaluation involves four articulated objects that look identical but differ in articulation: (a) a cabinet with a revolute joint on the left; (b) another cabinet with a revolute joint on the right; (c) a cabinet with a revolute joint at the bottom; and (d) a drawer with a prismatic joint. In scenarios (e)–(h), corresponding to objects (a)–(d), Tac-Man starts from the same position and in the same direction. The initial backward pull by the arm reveals variations in contact loss due to the differing articulations. Upon reaching a certain deviation threshold, the system triggers an adjustment in the arm’s pose. This adjustment, optimized by Eq. 23, aims to recover stable contact and re-align to the correct interaction direction. The arm then continues its backward pull. Through this iterative process of adjustment and execution, each object, starting from the same initial state, is maneuvered to a distinct final state reflective of its unique articulation. For video demonstrations, refer to Supplementary Video S1.
Figure 7: Manipulation under imperfect priors. Tac-Man adeptly addresses challenges posed by imperfect priors, which accurately identify the model type but contain erroneous parameters. (a) An example is an off-axis door, designed for space and energy efficiency, diverging from traditional models by not placing the hinge directly on one side. This variation can mislead prior assumptions regarding the arc trajectory’s radius. (b) We mirror this scenario by assigning the robot a pre-programmed arc trajectory to open a microwave oven with a revolute joint, incorporating a $10~{}\%$ radius error. This slight misalignment between the priors and actual parameters leads to a gradual loss, eventually causing the gripper to lose contact with the handle, resulting in a failed manipulation attempt. (c) In contrast, Tac-Man dynamically re-calibrates the arm’s pose to correct any unexpected errors, ensuring alignment with the correct interaction direction, going on to successfully open the microwave oven. This demonstration of adaptability underscores Tac-Man’s ability to complete tasks even with imperfect priors. For video demonstrations, refer to Supplementary Video S2.
@@
Tac-Man effectively demonstrates proficient manipulation of distinct objects in all four cases, as shown in Fig. 6(e)–(h), each corresponding to the configurations in Fig. 6(a)–(d). Starting from the same initial state with stable contact $\mathcal{C}_{1}$, Tac-Man employs an iterative two-stage cycle, beginning with an identical preliminary direction $R^{1}$. During the execution stage, tactile feedback is utilized to discern deviations from the correct interaction direction by observing changes in the contact points. For example, the contact in (e) exhibits a shear counterclockwise deviation, indicating a need for an additional clockwise rotation in the robot’s movement, and conversely for the shear clockwise deviation observed in (f). In the scenario depicted in (g), a deviation in the normal direction $\Delta N$ signals the requirement for an added rotation for recovery. Conversely, in (h), the contact remains stable as the arm’s movement is congruent with the correct direction. When a predetermined deviation threshold is reached, the system enters the recovery stage, employing the optimization outlined in Eq. 23 to compute the optimal arm pose that compensates for the observed deviation, thereby realigning the gripper to the correct interaction direction. Following this adjustment, the process resumes along the preliminary direction $R^{1}$. Through continual refinement to preserve stable contact, the arm successfully manipulates the objects into their intended states, as depicted in the last images in Fig. 6.
@@
Figure 8: Manipulation under unknown priors. Tac-Man showcases exceptional proficiency in managing scenarios characterized by unknown priors, particularly when object kinematics defy prior-based prediction or exceed the limits of current robotic perception. We illustrate this capability with two distinct examples. In scenario (a), the object’s articulation mechanism, which is concealed beneath a cover, involves motions that combine translation and rotation—a complexity not easily captured by standard prismatic or revolute joint models. Scenario (b) presents a handle executing a helical trajectory, with the precise determination of its pitch posing a substantial challenge. Through these examples, Tac-Man proves its adeptness at navigating the intricacies of complex kinematic patterns, underscoring its effectiveness in scenarios where traditional perception models may falter. For video demonstrations, refer to Supplementary Video S3.

### v3

@@ III-C Contact Representation
For the effective computation of Eq. 6 and optimization of Eq. 7, a suitable representation of $\mathcal{C}_{i}$ is crucial. To ensure compatibility with the homogeneous transformation matrix and to match the capabilities of current tactile sensors, we represent the contact $\mathcal{C}_{i}$ as a set of positions of all discretized points at the contact site, relative to the reference frame $\{i\}$. The position of each contact point within $\mathcal{C}_{i}$ is denoted by $\boldsymbol{p}^{i}$. To facilitate alignment with the homogeneous transformation matrix, the column vector representation of $\boldsymbol{p}^{i}$ is defined in an augmented format as follows:
$\boldsymbol{p}^{i}=\begin{bmatrix}p_{x}^{i}&p_{y}^{i}&p_{z}^{i}&1\end{bmatrix}^{\mathsf{T}},$ (8)
@@
Significantly, $\mathcal{C}_{0}$ includes all points on the gripper’s outer surface. Assuming the total number of points is $m$, $\mathcal{C}_{0}$ is represented as:
$\mathcal{C}_{0}=\{\boldsymbol{p}_{j}^{0}\mid j=1,\dots,m\}.$ (9)
For simplicity, we assume a flat gripper surface, a common assumption in current robotic grippers and easily achievable with tactile sensor fabrication. Consequently, $p_{x}^{0}$ is consistent across all points and set to 00, attainable through coordinate transformation. The position of a contact point at $\mathcal{C}_{i}$ is then characterized by the condition:
$|p^{i}_{1}|\geq\epsilon$ (10)
@@
Therefore, $\mathcal{C}_{i}$ for $i=1\dots n$ can be defined as:
$\mathcal{C}_{i}=\{\boldsymbol{p}^{i}\mid|p^{i}_{1}|\geq\epsilon\}.$ (11)
III-D Stable Contact Formulation
@@
The computation of stable contact involves addressing the gripper’s maximum elastic deformation, encompassing both normal and shear components. These are captured in the expressions for $f_{e}(\mathcal{C}_{i},\mathcal{C}_{0})$ and $\boldsymbol{e}$ within the context of stable contact. Ensuring that the contact deformation remains within the material’s elastic limits is crucial for maintaining a stable grip on the object:
$f_{e}(\mathcal{C}_{i},\mathcal{C}_{0})=\begin{bmatrix}\max\{|u_{1}|\mid\boldsymbol{u}\in\mathcal{C}_{i}\}\\ \max\left\{\left\|\begin{bmatrix}u_{2}-v_{2}\\ u_{3}-v_{3}\end{bmatrix}\right\|_{2}\;\middle|\;(\boldsymbol{u},\boldsymbol{v})\in\mathcal{K}_{0i}\right\}\end{bmatrix}$ (12)
@@
$\boldsymbol{e}=\begin{bmatrix}e_{n}&e_{s}\end{bmatrix}^{\mathsf{T}},$ (13)
where $d_{n}$ and $d_{s}$ denote the maximum normal and shear elastic deformation, respectively. For notation simplicity, we denote $|u_{1}|$ as $\Delta N$ and $\left\|\begin{bmatrix}u_{2}-v_{2}\\ u_{3}-v_{3}\end{bmatrix}\right\|_{2}$ as $\Delta S$ henceforth.
Similarly, we formulate $f_{s}(\mathcal{C}_{i},\mathcal{C}_{0})$ and $s$ to address concerns related to slipping in Eq. 2. Practically, to prevent sliding, the friction $f$ between the object and the gripper should not surpass the maximum static friction $f_{M}$. In our model, the friction at $\mathcal{C}_{i}$ is positively correlated with the shear deformation observed from $\mathcal{C}_{0}$:
@@
$\Delta S\leq\delta^{M}(f_{N}),\forall(\boldsymbol{u},\boldsymbol{v})\in\mathcal{K}_{0i},$ (15)
@@
$\delta^{M}(\Delta N)=\delta_{0}\Delta N,(\boldsymbol{u},\boldsymbol{v})\in\mathcal{K}_{0i},$ (16)
@@
$\delta_{0}=\frac{\delta^{M}(\Delta N_{0})}{\Delta N_{0}},(\boldsymbol{u},\boldsymbol{v})\in\mathcal{K}_{0i}.$ (17)
@@
$\frac{\Delta S}{\Delta N}\leq\delta_{0},\forall(\boldsymbol{u},\boldsymbol{v})\in\mathcal{K}_{0i},$ (18)
leading to the expression for $f_{s}(\mathcal{C}_{i},\mathcal{C}_{0})$:
$f_{s}(\mathcal{C}_{i},\mathcal{C}_{0})=\max\left\{\frac{\Delta S}{\Delta N}\;\middle|\;(\boldsymbol{u},\boldsymbol{v})\in\mathcal{K}_{0i}\right\},$ (19)
@@
$f_{d}(\mathcal{C}_{i},\mathcal{C}_{1})=\frac{\sum_{(\boldsymbol{u},\boldsymbol{v})\in\mathcal{K}_{1i}}\|\boldsymbol{u}-\boldsymbol{v}\|_{2}}{|\mathcal{K}_{1i}|}.$ (21)
@@ III-E Robot Pose Update
Having established the contact representation and computation methods for stable contact constraints, we should address the optimization in Eq. 7 to compute the value of $t$ for the determination of $\boldsymbol{T}_{r}^{i}$. However, in the absence of knowledge about the handle’s trajectory $J$, determining $t$ explicitly remains challenging. Therefore, we adopt an exploration mechanism by continuously executing with small increments in $t$ until any term in the constraints of optimization Eq. 7 reaches its upper bound, scaled by a safety margin $\alpha$ ($10<\alpha<$ 1). This process is concluded as
$\boldsymbol{T}_{r}=\boldsymbol{T}_{i}\boldsymbol{T}_{r}^{i},$ (22)
@@
To recover the contact to within the specified bounds after the execution stage, we employ the Kabsch algorithm [80] to compute the optimal transformation $\boldsymbol{T}^{r}_{i+1}$ for the recovery term in Eq. 6:
$\boldsymbol{T}^{r}_{i+1}=\operatorname*{arg\,min}_{\boldsymbol{T}^{r}_{i+1}\in SE(3)}\sum_{(\boldsymbol{u},\boldsymbol{v})\in\mathcal{K}_{1r}}\|\boldsymbol{T}^{r}_{i+1}\boldsymbol{u}-\boldsymbol{v}\|_{2}.$ (23)
@@
Figure 2: Real-world experiment setup. The experiments are conducted in a real-world environment using a Gen3 7DoF robotic arm paired with a Robotiq 2F-85 gripper. To equip it with the ability to sense contact geometry, we adopt two GelSight-type tactile sensors, each featuring a $7\times 7$ grid of labeled markers for point correspondence between frames.
@@ IV-A Experimental Setup
Figure 3: Visualization of the contact geometry detected by the sensors adopted. The GelSight-type sensors are adept at providing contact geometry information $\Delta N$ in the normal direction during object interactions. Labeled markers are used for accurate point correspondence to further enhance their effectiveness. Utilizing these features, $\Delta S$ is calculated and visualized using white arrows in the illustration. For better visibility, the magnitude of these arrows is exaggerated by a factor of $5$, ensuring a clearer visualization of the contact dynamics.
@@
Tac-Man is designed for broad compatibility with robotic systems capable of facilitating 6D movement in the gripper. In experiments, we utilized a Kinova Gen3 7DoF robotic arm equipped with a Robotiq 2F-85 gripper (Fig. 2). To assess the contact constraints as outlined in Sec. III, we integrated two GelSight-type tactile sensors into the gripper.
Tactile sensor
To further enhance our ability to validate these constraints, we adopt GelSight-type tactile sensors, which effectively provide geometry information at the contact surface. The back of the deformable pad is rigidly attached to a rigid acrylic sheet. Each sensor was specially fabricated with a unique $7\times 7$ grid of labeled markers (Fig. 3) on the front of the pad. Without any surrounding structural geometry, these markers share the same mechanical characteristics, allowing for calibration using consistent parameters. This grid aids in accurately identifying point correspondences between two contacts. Although this modification slightly reduces the resolution and depth reconstruction accuracy as compared to conventional GelSight sensors, its effectiveness in practical scenarios and in validating the contact constraints detailed in Sec. III is notable.
Figure 4: Selection concerns for the $\epsilon$ value. The compliance inherent in the tactile sensor material results in the activation of both the direct contact region and the adjacent transitional areas when contacting the handle. This characteristic underlines the need for a carefully adaptable $\epsilon$ value. Such adaptability ensures the precise identification of true contact points, thereby reducing the likelihood of detecting unreliable points.
@@
  • The threshold for identifying contact points, $\epsilon$ in Eq. 10, is set to be adaptable. This is due to the material’s flexibility, as shown in Fig. 4, which results in deformation not only in the contact region but also in the surrounding area upon contact with the object. This deformation relates to $\Delta N$ and the handle’s shape, among other factors. For adaptability, $\epsilon$ is defined as the minimal $\Delta N$ that ensures at least 8 contact points, exceeding the number of markers in a line (7 in our sensors), thus ensuring these points are non-colinear.
  • The normal elastic deformation limit, $e_{n}$ in Eq. 13, is set at $2~{}\mathrm{mm}$. This value is chosen based on observations that within this range, the system can revert to its initial state after releasing contact. It is important to note that this value represents a conservatively reduced normal elastic deformation for enforcing stricter constraints.
@@
  • The shear elastic deformation limit, $e_{s}$ in Eq. 13, is calculated as $e_{n}\delta_{0}$. Though smaller than the actual one, this value is deemed acceptable for stricter constraint enforcement.
  • The upper bound for another stable contact constraint, $d$ in Eq. 3, is experimentally determined to be $0.4~{}\mathrm{mm}$, ensuring it does not exceed $e_{n}$, $e_{s}$, and $\delta_{0}$ during manipulation.
  • The movement magnitude in the execution stage, the small increment in $t$ established in Eq. 7, is set as $\epsilon\delta_{0}$ to preserve stable contact during execution.
@@
hyper-parameter $\epsilon$ $e_{n}$ $e_{s}$ $\delta_{0}$ $d$ $\alpha$
value adaptive $2~{}\mathrm{mm}$ $3~{}\mathrm{mm}$ $1.5$ $0.4~{}\mathrm{mm}$ $0.6$
@@
Figure 5: $\boldsymbol{\delta_{0}}$ value acquisition. The experiment to obtain hyper-parameter $\delta_{0}$, as defined in Eq. 17: (a) The experiment begins with the arm pulling backward after the gripper has firmly grasped a slippery Lego piece. (b) Multiple trials are conducted at different values of $\Delta N$. Three representative trials are visualized to illustrate the process. The results follow the friction model well and $\delta^{M}$ can be easily identified. The results align well with the friction model, allowing for the straightforward identification of $\delta^{M}$. (c) The $\delta_{0}$ obtained from these multiple trials is visualized, and a value of $1.5$ is selected for safety considerations.
@@ IV-B Manipulation under Ambiguous Priors
Figure 6: Manipulation under ambiguous priors. Tac-Man showcases its proficiency in addressing the challenge caused by ambiguous priors, characterized by objects that, while visually indistinguishable, feature distinct articulation mechanisms. Our evaluation involves four articulated objects that look identical but differ in articulation: (a) a cabinet with a revolute joint on the left; (b) another cabinet with a revolute joint on the right; (c) a cabinet with a revolute joint at the bottom; and (d) a drawer with a prismatic joint. In scenarios (e)–(h), corresponding to objects (a)–(d), Tac-Man starts from the same position and in the same direction. The initial backward pull by the arm reveals variations in contact loss due to the differing articulations. Upon reaching a certain deviation threshold, the system triggers an adjustment in the arm’s pose. This adjustment, optimized by Eq. 23, aims to recover stable contact and re-align to the correct interaction direction. The arm then continues its backward pull. Through this iterative process of adjustment and execution, each object, starting from the same initial state, is maneuvered to a distinct final state reflective of its unique articulation. For video demonstrations, refer to Supp Video S1.
@@
Figure 7: Manipulation under imperfect priors. Tac-Man adeptly addresses challenges posed by imperfect priors, which accurately identify the model type but contain erroneous parameters. (a) An example is an off-axis door, designed for space and energy efficiency, diverging from traditional models by not placing the hinge directly on one side. This variation can mislead prior assumptions regarding the arc trajectory’s radius. (b) We mirror this scenario by assigning the robot a pre-programmed arc trajectory to open a microwave oven with a revolute joint, incorporating a $10~{}\%$ radius error. This slight misalignment between the priors and actual parameters leads to a gradual loss, eventually causing the gripper to lose contact with the handle, resulting in a failed manipulation attempt. (c) In contrast, Tac-Man dynamically re-calibrates the arm’s pose to correct any unexpected errors, ensuring alignment with the correct interaction direction, going on to successfully open the microwave oven. This demonstration of adaptability underscores Tac-Man’s ability to complete tasks even with imperfect priors. For video demonstrations, refer to Supp Video S2.
@@
Tac-Man effectively demonstrates proficient manipulation of distinct objects in all four cases, as shown in Fig. 6(e)–(h), each corresponding to the configurations in Fig. 6(a)–(d). Starting from the same initial state with stable contact $\mathcal{C}_{1}$, Tac-Man employs an iterative two-stage cycle, beginning with an identical preliminary direction $\boldsymbol{R}^{1}$. During the execution stage, tactile feedback is utilized to discern deviations from the correct interaction direction by observing changes in the contact points. For example, the contact in (e) exhibits a shear counterclockwise deviation, indicating a need for an additional clockwise rotation in the robot’s movement, and conversely for the shear clockwise deviation observed in (f). In the scenario depicted in (g), a deviation in the normal direction $\Delta N$ signals the requirement for an added rotation for recovery. Conversely, in (h), the contact remains stable as the arm’s movement is congruent with the correct direction. When a predetermined deviation threshold is reached, the system enters the recovery stage, employing the optimization outlined in Eq. 23 to compute the optimal arm pose that compensates for the observed deviation, thereby realigning the gripper to the correct interaction direction. Following this adjustment, the process resumes along the preliminary direction $\boldsymbol{R}^{i}$. Through continual refinement to preserve stable contact, the arm successfully manipulates the objects into their intended states, as depicted in the last images in Fig. 6.
@@
Figure 8: Manipulation under unknown priors. Tac-Man showcases exceptional proficiency in managing scenarios characterized by unknown priors, particularly when object kinematics defy prior-based prediction or exceed the limits of current robotic perception. We illustrate this capability with two distinct examples. In scenario (a), the object’s articulation mechanism, which is concealed beneath a cover, involves motions that combine translation and rotation—a complexity not easily captured by standard prismatic or revolute joint models. Scenario (b) presents a handle executing a helical trajectory, with the precise determination of its pitch posing a substantial challenge. Through these examples, Tac-Man proves its adeptness at navigating the intricacies of complex kinematic patterns, underscoring its effectiveness in scenarios where traditional perception models may falter. For video demonstrations, refer to Supp Video S3.